What will be the wavelength of ultrasonic waves in the air if the frequency of waves is 20khz
1
Expert's answer
2020-04-13T10:15:58-0400

The wavelength of ultrasonic waves in the air

"\lambda=\frac{\rm speed\; of\; sound}{\rm frequency}=\frac{c}{f}"

"\lambda=\frac{340\:\rm m/s}{20000\:\rm Hz}=0.017\:\rm m=17\: mm."